175, 180, 200, 212, 214, 215, 218,

225, 230, 235, 239, 245, 265, 290

Determine the Interquartile Range (IQR).
Answer #1

n = 14

Median is (14+1)/2th item =average of 7th & 8th items = (218+ 225)/2 = 221.5

Bottom half is given by:

175, 180, 200, 212, 214, 215, 218

The middle of these numbers = 212

So,

Q1 = First Quartile = 212

Upper half is given by:

225, 230, 235, 239, 245, 265, 290

The middle of these numbers = 239

So,

Q3 =Third Quartile = 239

So,

Interquartile Range = IQR = Q3 - Q1 = 239 - 212 = 27

So,

Answer is:

27